OpenStudy (anonymous):

Find S10 of 1 + 5 + 25 + 125 + 625 + 3125?

OpenStudy (anonymous):

take each number multiply it by 4 and add itself to it to get the next number 1*4 + 1 = 5 5*4 + 5 = 25 25*4 + 25 = 125 etc

OpenStudy (anonymous):

3125*4 + 3125 = 15625 (7) 15625*4 + 15625 = 78125 (8) 78125*4 + 78125 = 390625(9) 390625*4 + 390625 = YOUR ANSWER